# A Comprehensive Evaluation on Quantization Techniques for Large Language Models

### v1

@@ 4 Conclusions
In this paper, we have presented an experimental evaluation of Large language model quantization. We systematically decouple the commonly used quantization methods into two key steps: pre-quantization transformation and quantization error mitigation. After a comprehensive review of existing approaches, we choose scaling and rotation in pre-quantization transformation step, GPTQ and low-rank compensation in quantization error mitigation step to further conduct experimental evaluations on the same ground. We found that uses the optimized rotation and scaling, together with two compensation methods demonstrating the best performance. For quantization settings, asymmetric quantization and smaller group sizes yield better results. For MXFP4 quantization, we can adopt methods used in INT4 quantization. However, using scaling-based pre-quantization transformation demonstrates superior performance compared to INT4. To achieve optimal performance under MXFP4 quantization, further investigation is required to explore its full potential and limitations.

### v2

@@ 4 Conclusions
In this paper, we have presented an experimental evaluation of Large language model quantization. We systematically decouple the commonly used quantization into two key steps: pre-quantization transformation and quantization error mitigation. After a comprehensive review of existing approaches, we choose scaling and rotation in pre-quantization transformation step, GPTQ, and low-rank compensation in the quantization error mitigation step to further conduct experimental evaluations on the same ground. We found that using the optimized rotation and scaling, together with two compensation methods, demonstrates the best performance. For quantization settings, asymmetric quantization and smaller group sizes yield better results. And using asymmetric quantization on activations shows more benefits on performance compared with weights. For FP4 quantization, we can adopt methods used in INT4 quantization. However, rotation-based pre-quantization transformation yields no observable more performance gains for MXFP4 and NVFP4 than scaling-based methods. And we find that the format of scaling factors demonstrates a large effect. For FP8 format scaling factor, FP8 (E4M3) with an additional per-tensor FP32 scaling factor is better than FP8 (E8M0). For small group sizes, using a 4-bit scaling factor for weights is also feasible, albeit with some loss in accuracy. To achieve optimal performance under FP4 quantization, further investigation is required to explore FP4’s full potential and limitations.